CRY+VIS Interface Meeting
Date: 2021/7/30 13:00-13:50
Participants: Takahashi, Ikeda, Kimura, Sato, Ushiba, Tomaru, Aso, Yokozawa, Hirata, Miyo, Node
- IX: In finalization. Problem on cables (unbalance, coupling?).
- EX: In finalization. Cable (P7-3 pin#2) should be recovered.
- EY: The mirror was transported to Kamioka. FC was extended to the magnet. Regluing might be needed. Installation will be started in the next week.
- Corrugate tube fixing jig: Remained jigs were replaced in EY on 7/29.
- TM height setting: ITMY.
- Commissioning in air: on going in X-arm. Oplev calibration in IX will be finished on 8/2. That in EX will be done after the CAL alignment.
- Commissioning in vacuum:
- Overall commissioning:
- Acceptance check: Every Friday.
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers: GVs were implemented on 7/28-29.
- Operators to ALIGN the suspensions are necessary.
Participants: Takahashi, Hirata, Miyo, Sato, Fabian, Ishizaki, Sugiyama, Aso, Tanaka, Terrence
17598 Re-cabling Dsub-15 pin for BS oplev , takaaki.yokozawa
17573 Suspension status after M4 earthquake in Hida , kouseki.miyo
17582 Add [PICO] and [PICO OVERVIEW] menu to [sitemap]-[PICO] menu. , satoru.ikeda
17369 Added EPICS channel for VIS temperature adjustment , satoru.ikeda
17553 Comment , satoru.ikeda
17580 Alignment check of ITMX after the earthquake , takafumi.ushiba
17608 Setup of limit switches , ryutaro.takahashi
17481 ITMX MNH1 cable repairement , takafumi.ushiba
17589 Comment , takaaki.yokozawa
17587 Calibration of BF damper , ryutaro.takahashi
17602 Rebooter installation for pico motor drivers. (ITMY) , koji.nakagaki
17406 Setup of pre-isolator , ryutaro.takahashi
17552 Comment , ryutaro.takahashi
17584 Rebooter installation and network changes for pico motor drivers. (PRM) , koji.nakagaki
17549 Oplev hardware , fabian.arellano
17600 The oplev output viewport is dirty , fabian.arellano
17549 Oplev hardware , fabian.arellano
This is generated by python script in /users/Miyo/dropbox/Script/scrapping/main.py.
- Replaced the LVDT drivers for the BF damper in IX and EX.
- Repaired the cables in IX.
- Setup the limit switches in IX and EX.
- Checked the cable connections in IY and EX.
- Preparing for the accelerometers.
- We measured the height of the optical table and balanced it to make it horizontal.
- We used the laser level and measured the height at each suspension point with respect to the vertical centre of the flange.
- We redistributed the ballast mass on top to make it horizontal.
- The table is about 1 mm above the nominal value of 380 mm.
- The amount of ballast mass on top is 13.84 kg.
- We examined the cleanliness of the mirror with the new super ultra bright LED flashlight:
- It looks very clean!
- There a very few dust particles far away from the centre, so they're not a problem.
For the O3GK paper, I revisited the calculation of the effect of DAC and coil driver noise on DARM. Because no one measured the transfer function from individual coils to DARM we cannot propagate the noise directly. However, it is possibler to calculate upper limits using the transfer functions from L, P and Y of the mirrors.
- PR2: Hirata-san and Yano-san confirmed the optical table is 10 mm lower than nominal.
- PR3: They measured the height of the optical table.
- PR3: When testing Guardian, Miyo-kun shook the mirror accidentally and it moved -3 mrad in pitch. Fortunately, it was possible to recover it with the picomotor.
- Updated on Oplev.
- Type-A team: ACC preparation.
- Type-B team: Table height measurement in PRM/BS. OSEM offload in PR3.
IMC suspensions are ready in air  will finish 2021/04/21.
PRs are ready in air  will finish 2021/07/09.
ETMX is ready in air  will finish 2021/08/25.
ITMX is ready in air  will finish 2021/08/31.
ETMY is ready in air  will finish 2021/09/16.
ITMY is ready in air  will finish 2021/11/08.
BS is ready in air  will finish 2021/11/24.
SRs are ready in air  will finish 2021/12/20.
Output suspensions are ready in air  will finish 2021/12/21.
TMSY is ready in air  will finish 2021/??/??.
TMSX is ready in air  will finish 2021/??/??.
- Power leakage on 7/30.
The week from 8/2
- Takahashi: 8/2-6
- Sato: 8/2-6
- Hirata: 8/2-6
- Yano: 8/2-6
Ishizaki: 8/6-9 (Shift) -->canceled
The week from 8/9